∫(x² + 4xy) dx


1
Expert's answer
2020-04-13T19:13:41-0400

"\int (x^2 + 4xy) \, dx"

"= \int x^2 \, dx + \int 4xy \, dx \hspace{1 cm}" [ "\because \int (f(x) + g(x)) \, dx = \int f(x) \, dx + \int g(x) \, dx \,\," ( Where "f(x) \text{ and } g(x)" both are functions ) ]

"= \int x^2 \, dx + 4y \int x \, dx \hspace{1 cm}" [ "\because \int cf(x) \, dx = c \int f(x) \, dx \,\," ( Where "c" is a constant and "f(x)" is a function ) ]

"= \frac{x^3}{3} + 2yx^2 + C \hspace{1 cm}" [ "\because \int x^n \, dx = \frac{x^{n+1}}{n+1} + C \,\," ( Where "C" is a constant of integration) ]